OpenStudy (anonymous):

Simplify the ratio 4 : 22 : 30 Complete the pattern 486, 162, 54, 18

OpenStudy (anonymous):

4 : 22 : 30 2:11:15

OpenStudy (anonymous):

t1=486 t2=486/3=162 t3=162/3=54 t4=54/3=18 t5=18/3=6

OpenStudy (anonymous):

Got it?